Cliams on a gruop of people rise randomly and independently of each other through time at an average rate of 2 per month. Calculate the probability that 30 claims arise in a period of one year


1
Expert's answer
2022-03-14T17:46:20-0400

This is a Poisson distribution with "\mu=2*12=24."

"P(X=30)=e^{-24}\frac{24^{30}}{30!}=0.0363."